Operation Summary for October 7, 2011 - October 10, 2011 |
NuMI reported receiving 8.21E18 protons on target for the period from 10/3/11 to 10/10/11. Friday October 7 The day shift began with Operations monitoring the Pbar stack (47.1E10) and the protons in the Recycler (6.4E10), with the Main Injector (MI), Linac, and Booster operating normally, and with the NuMI and MiniBooNE experiments taking beam. During the day shift, the Neutron Therapy Facility reported at 8:10 AM that they were preparing their beam line for treating one patient later this morning. Operators reported at 8:54 AM that they were rebooting the MI-60 BPM system. Operations established beam to the Meson FTBF experiment T-1012 at 10:13 AM. Pbar personnel dumped the last remaining stack of antiprotons at 11:19 AM. Booster personnel began an aperture scan at 10:06 AM. Operators began preparing Pbar and their transfer lines for a 30-minute access to valve out a magnet with a LCW leak. At 1:33 PM, operators reported that a Switchyard (SY) corrector magnet power supply (S:V204) was tripping off on a high temperature indication; repairs on this power supply held off beam to T-992. Operators reported short downtimes due to problems with the NuMI BPMs, a MI RF vacuum pump (I:IPRF17), a Booster RF station (BRF17), and Linac quadrupole power supply (QF114). During the evening shift, Booster personnel ended their aperture scans at 4:26 PM. A Power Supply person reported at 4:49 PM that he had completed his work on S:V204; Operations resumed sending beam to T-992. At 8:35 PM, the V204 power supply tripped off again on a high temperature indication; a power supply person came in and began repairs. T-992 said they were done for the night due to the repairs and Operations disabled their beam. Saturday October 8 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, the I- Source high voltage power supply tripped off at 12:28 AM; it reset. At 1:56 AM, LRF3 tripped off for the first of two times in about three hours; it reset. At 3:13 AM, MIRF17 tripped off for the first of four times in about four hours; it reset. At 6:42 AM, MTA suffered a power outage; the main breaker for compressor building tripped off once again. Operators contacted the High Voltage Coordinator and the FESS Duty Electrician, who had power restored by 7:37 PM. The power tripped off again at 7:51 AM when the Cryo System Coordinator turned on the helium compressor. During the day shift, at 8:16 AM, a Safety System person reported a leak in the MiniBooNE target LCW system; Operations turned off beam to MiniBooNE and prepared the area for an access. FESS personnel restored power to MTA at 8:54 AM. Operations established beam to T-992 at 10:39 AM. Fluid Group personnel brought the regenerate CUB column #1 online at 1:51 PM. During the evening shift, Operators reported at 4:34 PM that the MiniBooNE water skid showed a reduce flow; an Instrumentation person came in to work on the system. MiniBooNE remained off. Operations disabled beam to T-992 at 10:25 PM. MIRF17 tripped off 10:36 PM; it reset. A Linac Klystron RF station (KRF3) tripped for the first of three times in six minutes; it reset. MIRF18 tripped off at 11:30 PM. Sunday October 9 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:05 AM, MIRF18 tripped off for the first of six times in about seven hours; it reset. At 12:32 AM, KRF3 tripped off for the first of seven times in about seven hours; it reset. At 1:53 PM, MIRF17 tripped off for the first of five times in about 20 minutes; operators bypassed the station after the last trip. At 3:14 AM, LRF5 tripped off for the first of six times in about 30 minutes; a Linac Technician came in after the last trip and brought the RF station back online at 4:10 AM. During the day shift, at 8:08 AM, KRF3 tripped off for the fist of many times throughout the shift; it reset. At 9:12 AM, the MI-40 lower power supply tripped off for the first of two times in about an hour; operators contacted a Power Supply person who said that there was diagnostics on the system to determine the problem. Operations established beam to T-992 at 10:15 AM. Operators adjusted a low flow switch for the MiniBooNE target LCW water skid at 2:41 PM and then established beam to MiniBooNE at 3:48 PM. During the evening shift, KRF3 tripped off at 5:14 PM. A Cryo System person informed Operations at 4:35 PM that the CHL control room would no longer be occupied at night. Operators reported a Booster East Gallery sump high water alarm at 6:09 PM; a Booster person investigated and said it would be okay until Monday. Operators bypassed the alarm. At 8:49 PM, operators reported continuing KRF3 problems, but that a Linac person was working on the station; she reported there were also problems with LRF5 that would have to wait until Monday morning for repairs. At 8:50 PM, BRF3 and BRF14 tripped off; they reset. Operations disabled beam to T-992 at 10 PM. Monday October 10 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:28 AM, BRF14 tripped off for the first of six times in five and a half hours; after the last trip, operators adjusted the anode program. At 4:18 AM, the MiniBooNE Horn power supply began tripping off on ground fault indications; operators contacted a MiniBooNE person who had Operations turn off the MiniBooNE beam until later in the morning. At 7:47 AM, Operations turned off the Linac beam so that Linac personnel could complete a phase calibration for the low energy RF stations. Operations Coordinator (OPSCO) The OPSCO reported that operators had found a tune for MI that significantly increased its efficiency over the weekend. |
|
More Information Comments and Suggestions |
| last modified 10/10/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |